# Haotong Yang    Zitong Wang    Shijia Kang    Siqi Yang    Wenkai Yu    Xu Niu    Yike Sun    Yi Hu    Zhouchen Lin    Muhan Zhang

### v1

@@ Abstract
While Large Language Models (LLMs) have demonstrated strong math reasoning abilities through Reinforcement Learning with Verifiable Rewards (RLVR), many advanced mathematical problems are proof-based, with no guaranteed way to determine the authenticity of a proof by simple answer matching. To enable automatic verification, a Reward Model (RM) capable of reliably evaluating full proof processes is required. In this work, we design a scalable data-construction pipeline that, with minimal human effort, leverages LLMs to generate a large quantity of high-quality “question-proof-check” triplet data. By systematically varying problem sources, generation methods, and model configurations, we create diverse problem-proof pairs spanning multiple difficulty levels, linguistic styles, and error types, subsequently filtered through hierarchical human review for label alignment. Utilizing these data, we train a proof-checking RM, incorporating additional process reward and token weight balance to stabilize the RL process. Our experiments validate the model’s scalability and strong performance from multiple perspectives, including reward accuracy, generalization ability and test-time guidance, providing important practical recipes and tools for strengthening LLM mathematical capabilities.
@@ 1 Introduction
Pretrained large language models (LLMs) have acquired the ability to produce complex Chain-of-Thought (CoT) through reinforcement learning-based post-training, enabling success on increasing difficult tasks, such as mathematical problems solving (DeepSeek, 2025; Huang and Yang, 2025). Reinforcement Learning relies on accurate and comprehensive reward signals to ensure stable and efficient training and to avoid reward hacking (Amodei et al., 2016). Even a small amount of imprecise rewards can impede learning (Gao et al., 2024; Xu et al., 2025) or steer model towards shortcut reasoning (Gao et al., 2022; Pang et al., 2023), substantially degrading reasoning performance. Meanwhile, RL involves evaluating highly diverse, on-policy data. A reward system must possess both generalizability and accuracy.
The classic Reinforcement Learning from Human Feedback (RLHF) pipeline collects human preference data and train a neural Reward Model (RM) to approximate human favor. While RLHF has been successful in scenarios such as chatbot and creative writing, it is constrained by the scope and quantity of human-annotations for complex reasoning, where expert labeling is extremely costly. According to Dekoninck et al. (2025), in Olympic-level math, annotating a single question requires domain experts (like IMO participants) to spend hours and costs hundreds of dollars. In addition, reasoning tasks often involve rich solution paths and vast search spaces, demanding more diverse supervision to achieve reward generalization.
@@
Unfortunately, in top-competition-level and research-level math, problems are predominantly proof-based. For example, in USAMO 2025  [^1] https://artofproblemsolving.com/wiki/index.php/2025_USAMO_Problems , 4 out of 6 problems are proof problem and the other 2 require both an answer and the proof. Considering that costly expert checking is now the only way to accurately evaluate these proof (Petrov et al., 2025), there is an increasingly urgent need to have a model that can verify the correctness of proofs. Because it determine the correctness of the proof and work as an RM, in the following part of the paper, we call the model to read the question (Q), check the proof (P) and return the check (C) results as “true/false” (T/F) as Proof RM.
@@
To scale the T/F annotations (Step 2 in Figure 1), we employ three LLM that judge each proof with a total of five times. The label will be accepted only under unanimous agreement. In view of human checking being expensive, to mitigate the misalignment between the unanimous agreement of LLM and human, we operate a combination-level sampling check to minimize human efforts. We partition data into different combinations by its data source, prompt method, and generating LLM, like “C1”, “C2” in Figure 1. For each combination , if sampled human judgments align strongly with LLMs, we treat that slice as silver-standard and keep them; otherwise we drop the entire combination, ensuring the overall consistency of the training data we collected. This pipeline yields 21k QPC training samples in about 100 hours and saving more than 90% [^2] The ratio is approximated by our one-by-one check for test set. See in Appendix A.2. time versus one-by-one check (see Section 3.2).
With sufficient QPC data, the labeling C constitutes a verifiable reward for training a proof RM, enabling an RLVR setup. In the RL stage, however, we observe that combining long reasoning process with binary T/F supervision introduce instability: drastic decrease in accuracy and increase in divergence, usually signaled by repetitive and syntactic incoherence. We hypothesize that this is due to the fact that noisy or speculative thinking process can still reach the correct verdict, thereby assigning incorrect reward to poor generations and inducing model collapse. We mitigate it with an “LLM-as-a-RM for RM” by supervising the thinking fluency. We also find the strong correlation between drastically varying output lengths and model collapse, which is solved by the introduction of a balanced token weight strategy combining sequence-level and token-level token weight. Our modified RL algorithm achieve scalable RL training with increasing performance on more than 312 steps with more than 18k samples and 112M trained tokens. Details can be found in Section 3.3.
@@
  • Analyzing the asymmetry of verification on math and pointing out the necessary of a proof RM;
@@ 2 Motivation: asymmetry of verification
2.1 Previous success based on strong asymmetry
LLMs have recently achieved breakthrough progress on mathematical problem solving. We first highlight that much of this progress can be attributed to a strong asymmetry in verification in current math training datasets and benchmarks. In Table 1, we summarize several important math benchmarks, along with the required knowledge of solving them and how to verifying their solutions. A striking pattern is that although solving these problems often requires mathematical knowledge ranging from middle-school level to highly creative advanced techniques, verifying correctness does not become comparably harder, typically can be automatically operated by codes or LMs. Moreover, in terms of reasoning steps, solving these tasks often requires multi-step, exploratory thinking, or branching into alternative solution paths. By contrast, the evaluation only focus on the output. It is precisely this asymmetry in verification that enables large-scale and highly efficient reinforcement learning, which in turn allows models to begin demonstrating strong capabilities on these tasks.
@@ 2.2 Limits of current asymmetry on math
Although RLVR is currently the most prevalent approach, its limitations are becoming increasingly apparent. First, the process of mathematics is even more important than only the conclusion—as argued by the mathematician William Thurston in his essay: “Rather, it reflected a continuing desire for human understanding of a proof, in addition to knowledge that the theorem is true.” (Thurston, 1994). The ideas, tools, lemmas derived to solve the problem matters more and highly develop math field. Yet vanilla RLVR provides no supervision of the reasoning process; its effectiveness implicitly rests on the assumption that “a correct result typically requires a correct derivation”. In contrast, a growing body of evidence shows that RLVR-trained models frequently exhibit correct-answer/incorrect-proof behavior: they may rely on incomplete induction, treat unproven conjectures as facts after checking only a handful of instances, or even invoke nonexistent theorems (Wen et al., 2025; Tarek and Beheshti, 2025; Petrov et al., 2025).
A more severe problem is that the asymmetry is not from the nature of math but from well-designing of the dataset creator. For more complicated math problem or reserach-level math problem, they are generally proof-based. Non-trivially verifiable results which cannot be easily guessed require significant expert ingenuity and effort. For example, FrontierMath (Glazer et al., 2025) requires 60 expert mathematicians (even one of contributor holds a Fields Medal) to create only 35 research-level problems with a non-trivially verifiable short output. Although this process can create useful benchmark, it is not scalable to construct training data for a reward model training.
2.3 Some attempt to verify math proof
@@
Several work try to provide “a verified answer” for the proof problems. MathConstruct (balunović2025mathconstructchallengingllmreasoning) extracts the constructive subproblem in the proof which can be verified by Python code. IneqMath (Sheng et al., 2025) translates the inequality proof into numerical bound estimation and relation prediction ($\geq$, $\leq$, >>, <<) tasks. DeepTheorem (Zhang et al., 2025b) translate the groundtruth theorem into statements either entailed by (T) or contradictory from (F) the original theorem and ask about the T/F of these induced statement. These methods indeed enhance the verification asymmetry of proof problems. However, such transformations substantially reduce the difficulty of the original tasks. In practice, these verification of results are highly susceptible to “shortcut learning”. As illustrated in Table 2, an undergraduate student with modest experience in math competitions can, with little effort, limited time and without constructing a complete proof, achieve a high probability of arriving at the correct answer merely by relying on prior knowledge of the problem’s structure. This observation suggests that such transformations are susceptible to hacking and lack appropriate defense mechanisms. Models learn from these supervision could also easily learn unreasonable thinking process, degrading the trustfulness and helpfulness of the models, motivating us to consider the direct evaluation of proofs themselves. The detail of the experiments are in Appendix A.3. Another related direction is formal language proofs, providing native verification methods at the expense of poor readability and logical redundancy. We leave them to Section 5 and discussion in Appendix A.1.
@@ 4.1 Reward accuracy
Table 4: Comparison with various baselines on testset, with the number in parantheses indicating the performance increase compared to the base models.
@@ 4.2 Generalization ability
Table 5: Comparison with various baselines on OPC test set (Dekoninck et al., 2025). Here we adopt the performance of closed-source LLM’s as is demonstrated in the original paper.
@@ A.4.5 Generated Negative sample analysis
A more obscure example of overlooking preconditions is contributed by the mask_replace method, with a combinatorics problem as follows. In this example, the generated solution overlooked the part on the construction of so called ”traps”, making the entire logic chain of the solution invalid. These kind of mistakes are also commonly seen in human-written solutions, either due to negligence or falsely beliefs that such preconditions are trivial.
@@
Additionally, another genre of negative samples generated by our pipeline that aligns with common human mistakes is the misuse of ”Without Loss of Generality” (WLOG). This type of mistake is ever-present in all levels of mathematical problem solving. The following is an example of such negative samples, also generated by the mask_replace method. The solution automatically converted the general polygon case to the quadrilateral case, which is neither explicitly stated nor obvious. The rationale for such a conversion can be obtained from the explanation excerpted from the reference solution. Without such reasoning, one is unable to distinguish between a valid solution and a mere analysis of a easy case.
@@ A.4.6 Details of human check
Table 7: Details of human check results. TP and TN stands for True Positive and True Negative respectively, equivalently cases where the results of LLM check align with human check results.
@@ A.5 Test set details
  • Test set: The generic test set, randomly split from our training set in the question level, but keep not only the QAs with consistent LLM judgment but also the ones with inconsistent judgment. The test set share a similar distribution of the problem source, extended methods and extending LLM but include different questions (as well as their proofs) to avoid data leakage. This test set mainly reflects the general performance improvement on diverse QA pairs by our Proof RM RL. This set contains 238 QA pairs from 51 questions, where 141 samples among them are positive and 97 samples are negative.
  • OPC: We adopt the testing split of (Dekoninck et al., 2025), consisting of 292 question proof pairs from 65 unique questions, where among them 185 are negative and 107 are positive. The problems are sourced from well-known regional and international mathematical contests, and all labeled by human experts.
  • Best-of-k test: To test whether our RM is helpful to guide Math Olympic solver in test-time scaling setting, we additionally collect new problems in 2024 and 2025 with LLM-generated proofs for each problem. These LLMs includes DeepSeek-v3.1, Gemini-2.5-flash, Gemini-2.5-pro, GPT-5, GPT-5-mini and Qwen3-Next-80B-A3B-thinking. We use different LLMs to provide sufficiently diverse proof and simulate the situation where the performance of the proofs are various. We also manually filtered some too easy or too hard problems, or some too short (naive negative) or truncated proofs. We finally collect 18 problems and generate 48 proofs for each problem, amounting the total number of question-proof pairs to 864.
@@ A.6 RL Hyper-parameters
Table 8: RL hyperparameters, uniform across all model size trainings.
@@ A.7 Model Collapse and LLM-as-an-RM-for-RM
In our RL training stage, model collapse occasional occurs where the accuracy suddenly drops and at the same time the output length also shows significant variant. We find just before the accuracy finally drop, there have some “aura” where the thinking process of the model has been seemed odd but it finally give a correct answer, as the example in the following.
@@
Therefore, there is error in proof.
@@
We can easily find some repeated and nonsense sentences (red parts), but finally the model provide a correct verdict (green parts), which means the reward will be positive and the repeated tokens will be further learned by models. Just after the aura appear with several steps, the model has finally collapse and it cannot give a final answer. Try to solve the problem, we need to find and rectify it once the aura appears.
Because this odd behavior is somehow easy to find, we just adopt an pretrained LLM to read the RM’s output. We call the model as RM-for-RM. Here, we provide only the generation (Check) of RM to the RM-for-RM instead of the input-generation pairs (QPC) to avoid the RM-for-RM Distracted by the math problem. Once the RM-for-RM find some odd behavior in the generation, whatever the T/F prediction is correct, the check will be set as incorrect check and the reward will be set to 0. This rectification can improve the training stability. We also use it as a more sensitive metric to avoid model collapse. If the odd generation ratio is with continuous five steps, we will stop the training and resume from the nearest checkpoint.
@@
The collapse we observed in our training is usually due to some incorrectly encourage introduced by these “surface-level” fault. We observe that in some steps, some checking generation will include some repeated phrases or words (like “so so so so”) but finally leads to a correct T/F judgment. If our RL process incorrectly encourage these samples, after several steps, the model may learn to generate repeated words more and more frequently, which finally lead to the collapse. So it is important to find these unexpected behaviors as early as possible and avoid encouraging them. In our training stage, we find at some steps, the odd generation occurred, and after 5-10 steps the entropy will significantly explode and the reward will drops to almost 0.

### v2

@@ Abstract
While Large Language Models (LLMs) have demonstrated strong math reasoning abilities through Reinforcement Learning with Verifiable Rewards (RLVR), many advanced mathematical problems are proof-based, with no guaranteed way to determine the authenticity of a proof by simple answer matching. To enable automatic verification, a Reward Model (RM) capable of reliably evaluating full proof processes is required. In this work, we design a scalable data-construction pipeline that, with minimal human effort, leverages LLMs to generate a large quantity of high-quality “question-proof-check” triplet data. By systematically varying problem sources, generation methods, and model configurations, we create diverse problem-proof pairs spanning multiple difficulty levels, linguistic styles, and error types, subsequently filtered through hierarchical human review for label alignment. Utilizing these data, we train a proof-checking RM, incorporating an “LLM-as-a-RM-for-RM” approach and balanced token weighting to stabilize the RL process. Our experiments validate the model’s scalability and strong performance from multiple perspectives, including reward accuracy, generalization ability and test-time guidance, providing important practical recipes and tools for strengthening LLM mathematical capabilities.
@@ 1 Introduction
Pretrained large language models (LLMs) have acquired the ability to produce complex Chain-of-Thought (CoT) through reinforcement learning-based post-training, enabling success on increasingly difficult tasks, such as mathematical problem solving (DeepSeek, 2025; Huang and Yang, 2025). Reinforcement Learning relies on accurate and comprehensive reward signals to ensure stable and efficient training and to avoid reward hacking (Amodei et al., 2016). Even a small amount of imprecise rewards can impede learning (Gao et al., 2024; Xu et al., 2025) or steer model towards shortcut reasoning (Gao et al., 2022; Pang et al., 2023), substantially degrading reasoning performance. Meanwhile, RL involves evaluating highly diverse, on-policy data. A reward system must possess both generalizability and accuracy.
The classic Reinforcement Learning from Human Feedback (RLHF) pipeline collects human preference data and trains a neural Reward Model (RM) to approximate human favor. While RLHF has been successful in scenarios such as chatbot and creative writing, it is constrained by the scope and quantity of human-annotations for complex reasoning, where expert labeling is extremely costly. According to Dekoninck et al. (2025), in Olympic-level math, annotating a single question requires domain experts (like IMO participants) to spend hours and cost hundreds of dollars. In addition, reasoning tasks often involve rich solution paths and vast search spaces, demanding more diverse supervision to achieve reward generalization.
@@
Unfortunately, in top-competition-level and research-level math, problems are predominantly proof-based. For example, in USAMO 2025  [^1] https://artofproblemsolving.com/wiki/index.php/2025_USAMO_Problems , 4 out of 6 problems are proof problems and the other 2 require both an answer and the proof. Considering that costly expert checking is now the only way to accurately evaluate these proofs (Petrov et al., 2025), there is an increasingly urgent need to have a model that can verify the correctness of proofs. Because it determines the correctness of the proof and work as an RM, in the following part of the paper, we call the model to read the question (Q), check the proof (P) and return the check (C) results as “true/false” (T/F) as Proof RM.
@@
To scale the T/F annotations (Step 2 in Figure 1), we employ three LLMs that judge each proof with a total of five times. The label will be accepted only under unanimous agreement. In view of human checking being expensive, to mitigate the misalignment between the unanimous agreement of LLM and human, we operate a combination-level sampling check to minimize human efforts. We partition data into different combinations by its data source, prompt method, and generating LLM, like “C1”, “C2” in Figure 1. For each combination, if sampled human judgments align strongly with LLMs, we treat that slice as silver-standard and keep them; otherwise we drop the entire combination, ensuring the overall consistency of the training data we collected. This pipeline yields 21k QPC training samples in about 100 hours and saving more than 90% [^2] The ratio is approximated by our one-by-one check for test set. See in Appendix A.2. time versus one-by-one check (see Section 3.2).
With sufficient QPC data, the labeling C constitutes a verifiable reward for training a proof RM, enabling an RLVR setup. In the RL stage, however, we observe that combining long reasoning process with binary T/F supervision introduces instability: drastic decrease in accuracy and increase in divergence, usually signaled by repetitive and syntactic incoherence. We hypothesize that this is due to the fact that noisy or speculative thinking process can still reach the correct verdict, thereby assigning incorrect reward to poor generations and inducing model collapse. We mitigate it with an “LLM-as-a-RM for RM” by supervising the thinking fluency. We also find the strong correlation between drastically varying output lengths and model collapse, which is solved by the introduction of a balanced token weight strategy combining sequence-level and token-level token weight. Our modified RL algorithm achieves scalable RL training with increasing performance on more than 312 steps with more than 18k samples and 112M trained tokens. Details can be found in Section 3.3.
@@
  • Analyzing the asymmetry of verification on math and pointing out the necessity of a proof RM;
@@ 2 Motivation: asymmetry of verification
2.1 Previous successes based on strong asymmetry
LLMs have recently achieved breakthrough progress on mathematical problem solving. We first highlight that much of this progress can be attributed to a strong asymmetry in verification in current math training datasets and benchmarks. In Table 1, we summarize several important math benchmarks, along with the required knowledge of solving them and how to verifying their solutions. A striking pattern is that although solving these problems often requires mathematical knowledge ranging from middle-school level to highly creative advanced techniques, verifying correctness does not become comparably harder, typically can be automatically operated by codes or LMs. Moreover, in terms of reasoning steps, solving these tasks often requires multi-step, exploratory thinking, or branching into alternative solution paths. By contrast, the evaluation only focuses on the output. It is precisely this asymmetry in verification that enables large-scale and highly efficient reinforcement learning, which in turn allows models to begin demonstrating strong capabilities on these tasks.
@@ 2.2 Limits of current asymmetry on math
Although RLVR is currently the most prevalent approach, its limitations are becoming increasingly apparent. First, the process of mathematics is even more important than only the conclusion—as argued by the mathematician William Thurston in his essay: “Rather, it reflected a continuing desire for human understanding of a proof, in addition to knowledge that the theorem is true.” (Thurston, 1994). The ideas, tools, lemmas derived to solve the problem matters more and highly develop the math field. Yet vanilla RLVR provides no supervision of the reasoning process; its effectiveness implicitly rests on the assumption that “a correct result typically requires a correct derivation”. In contrast, a growing body of evidence shows that RLVR-trained models frequently exhibit correct-answer/incorrect-proof behavior: they may rely on incomplete induction, treat unproven conjectures as facts after checking only a handful of instances, or even invoke nonexistent theorems (Wen et al., 2025; Tarek and Beheshti, 2025; Petrov et al., 2025).
A more severe problem is that the asymmetry is not from the nature of math but from well-designing of the dataset creator. For more complicated math problems or research-level math problems, they are generally proof-based. Non-trivially verifiable results which cannot be easily guessed require significant expert ingenuity and effort. For example, FrontierMath (Glazer et al., 2025) requires 60 expert mathematicians (even one of the contributors holds a Fields Medal) to create only 35 research-level problems with a non-trivially verifiable short output. Although this process can create a useful benchmark, it is not scalable to construct training data for a reward model training.
2.3 Attempts to verify math proofs
@@
Several works try to provide “a verified answer” for the proof problems. MathConstruct (balunović2025mathconstructchallengingllmreasoning) extracts the constructive subproblem in the proof which can be verified by Python code. IneqMath (Sheng et al., 2025) translates the inequality proof into numerical bound estimation and relation prediction ($\geq$, $\leq$, >>, <<) tasks. DeepTheorem (Zhang et al., 2025b) translate the groundtruth theorem into statements either entailed by (T) or contradictory from (F) the original theorem and ask about the T/F of these induced statement. These methods indeed enhance the verification asymmetry of proof problems. However, such transformations substantially reduce the difficulty of the original tasks. In practice, these verification of results are highly susceptible to “shortcut learning”. As illustrated in Table 2, an undergraduate student with modest experience in math competitions can, with little effort, limited time and without constructing a complete proof, achieve a high probability of arriving at the correct answer merely by relying on prior knowledge of the problem’s structure. This observation suggests that such transformations are susceptible to hacking and lack appropriate defense mechanisms. Models learn from these supervision could also easily learn unreasonable thinking process, degrading the trustfulness and helpfulness of the models, motivating us to consider the direct evaluation of proofs themselves. The detail of the experiments are in Appendix A.3. Another related direction is formal language proofs, providing native verification methods at the expense of poor readability and logical redundancy. We leave them to Section 5 and discussion in Appendix A.1.
@@ 4.1 Reward accuracy
Table 4: A comparison with various baselines on the test set, with the number in parentheses indicating the performance increase compared to the base models.
@@ 4.2 Generalization ability
Table 5: Comparison with various baselines on the OPC test set (Dekoninck et al., 2025). Here we adopt the performance of closed-source LLMs as demonstrated in the original paper.
@@ A.4.5 Generated Negative sample analysis
A more obscure example of overlooking preconditions is contributed by the mask_replace method, with a combinatorics problem as follows. In this example, the generated solution overlooked the part on the construction of so called ”traps”, making the entire logic chain of the solution invalid. These kinds of mistakes are also commonly seen in human-written solutions, either due to negligence or false beliefs that such preconditions are trivial.
@@
Additionally, another genre of negative samples generated by our pipeline that aligns with common human mistakes is the misuse of ”Without Loss of Generality” (WLOG). This type of mistake is ever-present in all levels of mathematical problem solving. The following is an example of such negative samples, also generated by the mask_replace method. The solution automatically converted the general polygon case to the quadrilateral case, which is neither explicitly stated nor obvious. The rationale for such a conversion can be obtained from the explanation excerpted from the reference solution. Without such reasoning, one is unable to distinguish between a valid solution and a mere analysis of an easy case.
@@ A.4.6 Details of human check
Table 7: Details of human check results. TP and TN stand for True Positive and True Negative respectively, equivalently cases where the results of the LLM check align with human check results.
@@ A.5 Test set details
  • Test set: The generic test set, randomly split from our training set in the question level, but keep not only the QAs with consistent LLM judgment but also the ones with inconsistent judgment. The test set shares a similar distribution of the problem source, extended methods and extending LLM but includes different questions (as well as their proofs) to avoid data leakage. This test set mainly reflects the general performance improvement on diverse QA pairs by our Proof RM RL. This set contains 238 QA pairs from 51 questions, where 141 samples among them are positive and 97 samples are negative.
  • OPC: We adopt the testing split of (Dekoninck et al., 2025), consisting of 292 question proof pairs from 65 unique questions, where among them 185 are negative and 107 are positive. The problems are sourced from well-known regional and international mathematical contests, and all are labeled by human experts.
  • Best-of-k test: To test whether our RM is helpful to guide Math Olympic solver in test-time scaling setting, we additionally collect new problems in 2024 and 2025 with LLM-generated proofs for each problem. These LLMs include DeepSeek-v3.1, Gemini-2.5-flash, Gemini-2.5-pro, GPT-5, GPT-5-mini and Qwen3-Next-80B-A3B-thinking. We use different LLMs to provide sufficiently diverse proof and simulate the situation where the performance of the proofs is various. We also manually filtered some too easy or too hard problems, or some too short (naive negative) or truncated proofs. We finally collect 18 problems and generate 48 proofs for each problem, amounting the total number of question-proof pairs to 864.
@@ A.6 RL Hyper-parameters
Table 8: The RL hyperparameters, uniform across all model size trainings.
@@ A.7 Model Collapse and LLM-as-an-RM-for-RM
In our RL training stage, model collapse occasionally occurs where the accuracy suddenly drops and at the same time the output length also shows significant variation. We find just before the accuracy finally drops, there is some “aura” where the thinking process of the model has been seemed odd but it finally gives a correct answer, as the example in the following.
@@
Therefore, there is an error in the proof.
@@
We can easily find some repeated and nonsense sentences (red parts), but finally the model provide a correct verdict (green parts), which means the reward will be positive and the repeated tokens will be further learned by models. Just after the aura appear with several steps, the model has finally collapsed and it cannot give a final answer. Try to solve the problem, we need to find and rectify it once the aura appears.
Because this odd behavior is somehow easy to find, we just adopt an pretrained LLM to read the RM’s output. We call the model as RM-for-RM. Here, we provide only the generation (Check) of RM to the RM-for-RM instead of the input-generation pairs (QPC) to avoid the RM-for-RM being distracted by the math problem. Once the RM-for-RM finds some odd behavior in the generation, whatever the T/F prediction is correct, the check will be set as an incorrect check and the reward will be set to 0. This rectification can improve the training stability. We also use it as a more sensitive metric to avoid model collapse. If the odd generation ratio is with a continuous five steps, we will stop the training and resume from the nearest checkpoint.
@@
The collapse we observed in our training is usually due to some incorrectly encourage introduced by these “surface-level” faults. We observe that in some steps, some checking generation will include some repeated phrases or words (like “so so so so”) but finally leads to a correct T/F judgment. If our RL process incorrectly encourages these samples, after several steps, the model may learn to generate repeated words more and more frequently, which finally leads to the collapse. So it is important to find these unexpected behaviors as early as possible and avoid encouraging them. In our training stage, we found at some steps, the odd generation occurred, and after 5-10 steps the entropy will significantly explode and the reward will drop to almost 0.